OpenStudy (anonymous):

The measure of a pair of vertical angles equal (8x-24) raised to 0 degree and (6x) raised to the 0 degree. What does the measure of each angle equal?

OpenStudy (mertsj):

Use this definition: \[a^0=1\]

OpenStudy (anonymous):

so do i set it ip like (8x-24) =1

OpenStudy (mertsj):

Read your problem carefully. Do you know what vertical angles are?

OpenStudy (anonymous):

opposite of eac other but share the same vertex?

OpenStudy (mertsj):



OpenStudy (mertsj):

Those are vertical angles. Do you know that vertical angles are congruent?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

You problem says: "the measure of a pair of vertical angles" So right off we know that they will have the same measure, right?

OpenStudy (anonymous):

right

OpenStudy (mertsj):

And then it says that the first one is : \[(8x-24)^0\]

OpenStudy (mertsj):

Isn't that what it says?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

And remember, I just posted the definition that: \[a^0=1\]

OpenStudy (mertsj):

So if we replace "a" with (8x-24) , what do we get for: \[(8x-24)^0\]

OpenStudy (anonymous):

1

OpenStudy (mertsj):

Yes and since vertical angles are equal both angles are equal to 1